### v2

# Mixed-Integer Optimal Control via Reinforcement Learning: A Case Study on Hybrid Vehicle Energy Management

Abstract
Many optimal control problems require the simultaneous output of continuous and discrete control variables. Such problems are usually formulated as mixed-integer optimal control (MIOC) problems, which are challenging to solve due to the complexity of the solution space. Numerical methods such as branch-and-bound are computationally expensive and unsuitable for real-time control. This brief proposes a novel continuous-discrete reinforcement learning (CDRL) algorithm, twin delayed deep deterministic actor-Q (TD3AQ), for MIOC problems. TD3AQ combines the advantages of both actor-critic and Q-learning methods, and can handle the continuous and discrete action spaces simultaneously. The proposed algorithm is evaluated on a plug-in hybrid electric vehicle (PHEV) energy management problem, where real-time control of the continuous variable, engine torque, and discrete variables, gear shift and clutch engagement/disengagement, is essential to maximize fuel economy while satisfying driving constraints. Simulation results on different drive cycles show that TD3AQ achieves near-optimal control compared to dynamic programming (DP) and outperforms baseline reinforcement learning algorithms.
Index Terms:
Mixed-integer optimal control, reinforcement learning, hybrid action space, hybrid electric vehicle energy management.
I Introduction
MIXED-INTEGER optimal control (MIOC) problems commonly arise in various engineering applications, such as assignment problems [1], control of piecewise-affine (PWA) systems [2] and hybrid systems [3]. These problems are particularly challenging as they require optimization of a system that involves both continuous and discrete control variables.
One common approach for solving MIOC problems is to formulate them as mixed-integer programming (MIP) models and apply the branch-and-bound method [4]. This method involves dividing the solution space into smaller branches and using a bounding function to determine if each branch can be eliminated. Although it can provide globally optimal solutions, it is computationally expensive and time-consuming. Another approach is to use dynamic programming (DP), which breaks the problem into smaller sub-problems and solves them recursively [5]. However, DP requires discretization of continuous variables and knowledge of the entire state space in advance, making it less suitable for real-world optimal control. Heuristic algorithms offer feasible and fast solutions by leveraging problem-specific knowledge to guide the search [6, 7], but they may not provide optimality guarantees and require customization for each problem structure [7].
The development of reinforcement learning (RL) has led to a new paradigm for solving complex control problems [8, 9]. RL learns a policy that maps the current state of the system to an action that maximizes the expected cumulative reward. With the learned policy, RL can reach a near-optimal solution without requiring any prior knowledge of the environment [10]. However, existing RL algorithms are mostly suitable for continuous or discrete action spaces, which cannot be directly applicable to MIOC problems.
Discretizing continuous variables into finite discrete values and applying discrete RL methods, such as Rainbow [11], is a straightforward approach to tackle MIOC problems. However, this approach becomes impractical for high-dimensional control problems as it necessitates a substantial number of discrete values to achieve accurate outcomes.
In recent years, a substantial amount of RL research has focused on a special type of MIOC problems called parameterized action Markov decision process (PAMDP) [12]. The key of PAMDP is to formulate the problem as a hierarchical structure where each discrete action is associated with a set of continuous parameters. One approach is to alternate between optimizing the discrete action selection with Q-learning and the parameter selection with a policy search method [12]. Another approach called parameterized deep deterministic policy gradient (PA-DDPG) is to directly output the continuous parameters and weights for discrete actions from DDPG [13] and select the discrete action with the highest weight [14]. Parameterized deep Q-networks (P-DQN) combine the actor-critic architecture and DQN, where the actor network first outputs parameters for all discrete actions, and then the Q-network selects the action with the highest Q-value [15]. Other approaches utilize neural ordinary differential equations as state-value networks [16] or conditional variational auto-encoders [17] to handle the parameterized action space. Alternatively, some methods handle the mixed-integer action space in a unified manner. These approaches add extra layers [18] or parallel networks [19] to generate continuous and discrete actions, and then optimize the policy using techniques such as dueling DQN or proximal policy optimization.
In the domain of mixed-integer control for hybrid electric vehicle (HEV) energy management, previous studies have employed various approaches, including utilizing frameworks similar to PA-DDPG [20] or P-DQN [21]. Tang et al. [22] combined DQN and DDPG, obtaining the gear ratio and engine throttle independently. However, no studies used RL to concurrently tackle the optimization of gear shifting and engine on/off control, which represents an unexplored research area in this field.
Inspired by the aforementioned research, this brief proposes a novel continuous-discrete reinforcement learning (CDRL) algorithm called twin delayed deep deterministic actor-Q (TD3AQ) for MIOC problems. The algorithm is applied to address an energy management problem for a plug-in hybrid electric vehicle (PHEV), aiming to demonstrate its effectiveness. The main contributions of this study are summarized as follows:
  1. 1)
    A state-of-the-art CDRL algorithm, TD3AQ, is proposed to address the mixed-integer action space. TD3AQ combines the advantages of both actor-critic and Q-learning methods, providing a unified solution for MIOC problems. Comparative evaluations demonstrate that TD3AQ achieves near-optimal solutions compared to dynamic programming (DP) and outperforms previous baselines [14, 23, 11]. Moreover, the sub-millisecond computation time of TD3AQ renders it suitable for real-time control applications.
  2. 2)
    This work pioneers the simultaneous optimization of two discrete variables, gear shift and clutch engagement/disengagement, using reinforcement learning for PHEV energy management. In contrast to previous studies [22] that rely on rule-based methods, this research emphasizes the systematic optimization [24] of gear shift and clutch engagement/disengagement for further fuel consumption reduction. The proposed approach is expected to yield superior energy management performance.
II Continuous-Discrete Reinforcement Learning
II-A Reinforcement Learning
Reinforcement learning is a learning approach for optimal control by interacting with the environment. The objective is to develop an agent that can learn to make optimal decisions by maximizing a cumulative reward signal. Typically, this problem is formulated as a Markov decision process (MDP), which consists of a state space $\mathcal{S}$, an action space $\mathcal{A}$, an initial state distribution $p_{0}$, and a transition probability $p(s^{\prime},r|s,a)$ that specifies the probability of transitioning to state $s^{\prime}$ and receiving reward $r$ when taking action $a$ in state $s$, where $r(s,a)\in\mathbb{R}$. The agent’s goal is to maximize the expected cumulative reward $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, where $\gamma\in[0,1]$ is the discount factor.
According to the Bellman optimality principle, the value of a state under an optimal policy is the maximum expected return achievable from that state. Reinforcement learning algorithms gradually refine value function estimates and converge to the optimal policy, which maximizes cumulative reward.
To achieve this goal, there are two mainstream algorithms: value-based methods and policy-based methods. Value-based methods learn the value function $V(s)$ or the action-value function $Q(s,a)$, which represent the expected cumulative reward starting from state $s$ or taking action $a$ in state $s$, respectively. The optimal policy $\pi^{*}(s)$ can be derived from the value function $V(s)$ or the action-value function $Q(s,a)$ by selecting the action that maximizes $Q(s,a)$, i.e., $\pi^{*}(s)=\arg\max_{a}Q(s,a)$. Policy-based methods, on the other hand, learn the parameterized policy $\pi_{\theta}(s)$ that can select actions without consulting a value function. The policy is updated by gradient ascent on a performance measure $J(\theta)$, where $\theta$ represents the parameters of the policy network. The policy network is a function approximator that maps the state $s$ to the action $a$.
Actor-critic methods combine value-based and policy-based approaches. They learn both an approximate value function and a parameterized policy simultaneously, enabling efficient learning in complex environments.
II-B Mixed-Integer Action Space
In this study, we explore the mixed-integer action space from the perspective of a parallel architecture, where discrete and continuous actions are independent of each other. For instance, in a PHEV energy management problem, the gear shift command could be the discrete action, while the engine torque could be the continuous action. These actions can be executed simultaneously and separately. Such mixed-integer action spaces can be defined as follows:
$\mathcal{A}=\{a=(a^{d},a^{c})\mid a^{d}\in\mathcal{A}_{d},a^{c}\in\mathcal{A}_{c}\}$ (1)
where $\mathcal{A}_{d}\subseteq\mathbb{N}^{m}$ represents a finite set of discrete action spaces, $\mathcal{A}_{c}\subseteq\mathbb{R}^{n}$ denotes a continuous action space, and $a^{d}$ and $a^{c}$ are the discrete and continuous actions, respectively. The dimensionality of the action space is determined by $m+n$. To handle the action space for multi-dimensional discrete variables, a composite action vector can be employed. This vector has a length of $d_{1}\times d_{2}\times\cdots\times d_{m}$, where $d_{i}$ represents the number of discrete actions for the $i$-th dimension. Each element of the vector corresponds to a distinct combination of discrete actions.
II-C TD3AQ Algorithm
Figure 1: The forward propagation process of TD3AQ. The continuous action $a^{c}$ is generated by the actor network $\mu_{\theta}$, the subscript of $a^{c}$ represents the dimension of the action space, and the discrete action $a^{d}$ is selected by $\arg\max$ operation on the Q-values.
To address the above mixed-integer action space, we propose a novel model-free algorithm, TD3AQ, which is based on the actor-critic framework. The forward propagation process of TD3AQ is shown in Fig. 1. The actor network $\mu_{\theta}$ is used to generate the continuous actions $a^{c}$, then $a^{c}$ and state $s$ are fed into the Q-network to obtain the Q-values $Q_{1},Q_{2},\ldots,Q_{k}$ for each discrete action. Finally, the discrete action with the highest Q-value is selected as $a^{d}$. The details of the TD3AQ algorithm are elaborated below.
II-C1 Deep Deterministic Actor-Q
Considering the vanilla DDPG algorithm [13], the Bellman equation is given by
$Q_{\omega}(s,a)=\mathbb{E}_{r,s^{\prime}}\left[r+\gamma Q_{\omega}\left(s^{\prime},\mu_{\theta}(s^{\prime})\right)\right]$ (2)
where $\omega$ represents the parameters of the Q-network, $\mu_{\theta}$ represents the actor network that approximates the optimal action in the next state $s^{\prime}$, and $\gamma$ is the discount factor. The loss function to satisfy the Bellman equation is defined as
$L(\omega)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{B}}\left[\left(y-Q_{\omega}(s,a)\right)^{2}\right]$ (3)
where $\mathcal{B}$ is the replay buffer, a minibatch is sampled from $\mathcal{B}$ to update the Q-network, $y$ is the target value, which is defined as
$y=r+\gamma Q_{\omega^{\prime}}\left(s^{\prime},\mu_{\theta^{\prime}}(s^{\prime})\right)$ (4)
where $\omega^{\prime}$ and $\theta^{\prime}$ are the target networks to stabilize the training process. It is worth noting that the Eq. (2) holds only when the actor network $\mu_{\theta}$ can provide the action that maximizes $Q_{\omega}(s,a)$. Assuming $Q_{\omega}$ is differentiable with respect to $a$, the actor network can be updated using the gradient ascent method by maximizing
$J(\theta)=\mathbb{E}_{s\sim\mathcal{B}}\left[Q_{\omega}(s,\mu_{\theta}(s))\right]$ (5)
Note that the parameters of $Q_{\omega}$ are fixed during the actor network update.
Deep deterministic actor-Q (DDAQ) [25] adopts the same architecture as DDPG, but instead of approximating a single Q-value, it approximates $k$ Q-values for the $k$ discrete actions. Therefore, the Bellman equation can be rewritten as
$Q_{\omega}(s,a)=\mathbb{E}_{r,s^{\prime}}\left[r+\gamma\max_{a^{d}\in\mathcal{A}_{d}}Q_{\omega}\left(s^{\prime},a^{d},\mu_{\theta}(s^{\prime})\right)\right]$ (6)
Additionally, the loss function can be defined as
$L(\omega)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{B}}\left[\left(y-Q_{\omega}(s,a^{d},a^{c})\right)^{2}\right]$ (7)
where
$y=r+\gamma\max_{a^{d}\in\mathcal{A}_{d}}Q_{\omega^{\prime}}\left(s^{\prime},a^{d},\mu_{\theta^{\prime}}(s^{\prime})\right)$ (8)
When updating the actor network, since the discrete action with the highest Q-value is selected as the output, the objective function is defined as
$J(\theta)=\mathbb{E}_{s\sim\mathcal{B}}\left[\max_{a^{d}\in\mathcal{A}_{d}}Q\left(s,a^{d},\mu_{\theta}(s)\right)\right]$ (9)
where the maximum Q-value is used as the objective function. In previous work (P-DQN) [15], the objective function was defined as the sum of all Q-values, which posed a problem when back propagating gradients: updating the weights in the Q-network during back propagation affected all weights in the network, including those unrelated to the action with the highest Q-value, potentially causing the algorithm to update in an incorrect direction. DDAQ addresses this issue by using the maximum Q-value as the objective function, ensuring that only the weights corresponding to the action $a^{d}$ with the highest Q-value are updated.
II-C2 TD3AQ
DDPG has shown potential in achieving good performance, but its robustness is limited due to its sensitivity to hyperparameters. This is mainly because DDPG tends to overestimate Q-values, which results in poor policy updates. To address this issue, we combine DDAQ and TD3 [26] to form TD3AQ. Three key modifications to improve the performance are introduced.
Target policy smoothing
a noise sampled from a normal distribution is added to the action generated by the target actor network. This technique acts as a regularizer that prevents the actor network from converging to a suboptimal solution. Specifically, the action was smoothed by adding a clipped noise, defined as follows:
$\tilde{a^{c}}=\operatorname{clip}(\mu_{\theta^{\prime}}(s^{\prime})+noise,a_{min}^{c},a_{max}^{c})$ (10)
where $a_{min}^{c}$ and $a_{max}^{c}$ are the lower and upper bounds of the continuous action, respectively. The noise is defined as:
$noise=\operatorname{clip}(\epsilon,\epsilon_{l},\epsilon_{h}),\quad\epsilon\sim\mathcal{N}(0,\sigma^{2})$ (11)
where $\epsilon_{l}$ and $\epsilon_{h}$ are the lower and upper bounds of the noise $\epsilon$, respectively. The noise $\epsilon$ is sampled from a normal distribution with mean 0 and standard deviation $\sigma$.
Clipped double Q-learning
to mitigate the overestimation of Q-values, a twin Q-network is introduced to DDAQ. When updating the Q-network, the minimum of the two selected Q-values is used as the target value:
$y=r+\gamma\min_{j=1,2}\left(\max_{a^{d}\in\mathcal{A}_{d}}Q_{\omega_{j}^{\prime}}\left(s^{\prime},a^{d},\tilde{a^{c}}\right)\right)$ (12)
then both Q-networks are updated by minimizing the mean squared error between the target value and the Q-value:
$L(\omega_{j})=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{B}}\left[\left(y-Q_{\omega_{j}}(s,a^{d},a^{c})\right)^{2}\right],\quad j=1,2$ (13)
Delayed updates
the actor network and target networks are updated less frequently than the Q-networks. This helps to stabilize the training process since slowing down the actor allows the critic to learn more accurate Q-values before letting it guide the actor. The networks are updated every two iterations as recommended in [26].
The proposed TD3AQ pseudo-code is shown in Algorithm 1, with all the proposed improvements combined.
Input: Gradient stepsizes $\{\alpha,\beta,\rho\}\geq 0$, exploration noise $\eta=\mathcal{N}(0,\delta^{2})$, exploration parameter $\varepsilon$, minibatch size $B$, empty replay buffer $\mathcal{B}$, target smooth noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$, target smooth noise bounds $\epsilon_{l}$, $\epsilon_{h}$, discount factor $\gamma$, policy update frequency $K$.
1 Initialize Q-network $Q_{\omega_{1}}$, $Q_{\omega_{2}}$, and actor network $\mu_{\theta}$ with random weights $\omega_{1}$, $\omega_{2}$, $\theta$;
2 Initialize target networks $\omega_{1}^{\prime}\leftarrow\omega_{1}$, $\omega_{2}^{\prime}\leftarrow\omega_{2}$, $\theta^{\prime}\leftarrow\theta$;
3 Initialize replay buffer $\mathcal{B}$;
4 for $t=1,2,\ldots,T$ do
5       Clear local gradients $\mathrm{d}\omega_{1}\leftarrow 0$, $\mathrm{d}\omega_{2}\leftarrow 0$, $\mathrm{d}\theta\leftarrow 0$;
6       Compute continuous action with exploration noise $a_{t}^{c}=\operatorname{clip}\left(\mu_{\theta}\left(s_{t}\right)+\eta,a_{min}^{c},a_{max}^{c}\right)$;
7       Compute discrete action $a_{t}^{d}$ according to $\varepsilon$-greedy
$\begin{cases}&\mathrm{Pr}(a_{t}^{d}=\text{a sample randomly from }\mathcal{A}_{d})=\varepsilon,\\ &\mathrm{Pr}(a_{t}^{d}=\arg\max_{a^{d}}Q_{\omega_{1}}\left(s_{t},a^{d},a_{t}^{c}\right))=1-\varepsilon\end{cases}$
Take action $a_{t}=(a_{t}^{d},a_{t}^{c})$, observe reward $r_{t}$ and the next state $s_{t+1}$;
8       Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in replay buffer $\mathcal{B}$;
9       Sample $B$ transitions $\{(s_{i},a_{i},r_{i},s_{i+1})\}_{i\in[B]}$ from $\mathcal{B}$;
10       Compute target continuous action
$\tilde{a_{i}^{c}}=\operatorname{clip}(\mu_{\theta^{\prime}}(s_{i+1})+\operatorname{clip}(\epsilon,\epsilon_{l},\epsilon_{h}),a_{min}^{c},a_{max}^{c})$
Compute the target value
$y_{i}=r_{i}+\gamma\min_{j=1,2}\left(\max_{a^{d}\in\mathcal{A}_{d}}Q_{\omega_{j}^{\prime}}\left(s_{i+1},a^{d},\tilde{a_{i}^{c}}\right)\right)$
Update the two Q-networks ($j=1,2$) by one step of gradient descent
$\omega_{j}\leftarrow\omega_{j}-\alpha\nabla_{\omega_{j}}\frac{1}{|B|}\sum_{i\in[B]}\left(Q_{\omega_{j}}\left(s_{i},a_{i}^{d},a_{i}^{c}\right)-y_{i}\right)^{2}$
if t mod $K$ == 0 then
11             Update the actor network by one step of gradient ascent
$\theta\leftarrow\theta+\beta\nabla_{\theta}\frac{1}{|B|}\sum_{i\in[B]}\max_{a^{d}}Q_{\omega_{1}}\left(s_{i},a^{d},\mu_{\theta}\right)$
Update the target networks ($j=1,2$) with
$\omega_{j}^{\prime}\leftarrow\rho\omega_{j}+(1-\rho)\omega_{j}^{\prime},\quad\theta^{\prime}\leftarrow\rho\theta+(1-\rho)\theta^{\prime}$
12       end if
13
14 end for
Algorithm 1 Pseudo-code of the TD3AQ Algorithm
III PHEV Energy Management Problem Formulation
To validate the proposed TD3AQ algorithm, this work considers a PHEV energy management problem, which presents a challenging MIOC problem as it requires simultaneous control of continuous power distribution and discrete gear selection for a highly non-linear hybrid system. The section begins with an introduction to the PHEV powertrain modeling, followed by the MIOC problem formulation.
III-A PHEV Powertrain Configuration
Figure 2: The parallel hybrid powertrain architecture.
This study focuses on a parallel hybrid electric vehicle [27]. The powertrain architecture comprises an internal combustion engine, an electric motor, a clutch, and an automated manual transmission, as depicted in Fig. 2. The vehicle’s HEV mode can be altered by engaging or disengaging the clutch between the engine and the motor, providing three modes of operation: motor-only, engine-only, and hybrid-drive.
III-B Vehicle Dynamics
The vehicle dynamics are described by the following equations:
$T_{w}=\left[ma+\frac{1}{2}C_{d}\rho Av^{2}+\mu mg\cos\theta+mg\sin\theta\right]r$ (14)
$\omega=\frac{v}{r}i_{d}i_{g}$ (15)
where $T_{w}$ is the wheel torque, $m$ is the vehicle mass, $a$ is the vehicle acceleration, $C_{d}$ is the air drag coefficient, $\rho$ is the air density, $A$ is the vehicle cross section area, $\omega$ is the shaft angular velocity, $v$ is the vehicle velocity, $\mu$ is the rolling resistance coefficient, $g$ is the gravitational acceleration, $\theta$ is the road grade, $r$ is the tyre radius, $i_{d}$ is the final drive gear ratio, and $i_{g}$ is the transmission gear ratio.
The torque on the wheel can be denoted as:
$T_{w}=(cT_{e,d}+T_{m})i_{d}i_{g}\eta_{d}\eta_{g}-T_{b}$ (16)
where $T_{e,d}$ and $T_{m}$ represent the driving torque transmitted from the engine and motor to the shaft, respectively. $c$ is the clutch state with $c=1$ representing the clutch engaged and $c=0$ representing the clutch disengaged. $\eta_{d}$ is the final drive efficiency, $\eta_{g}$ is the transmission efficiency, and $T_{b}$ is the mechanical brake torque.
III-C Engine Model
Figure 3: The engine fuel rate map, where the boundary represents the maximum engine torque.
The engine torque is separated into two parts: the driving torque $T_{e,d}$ and the idle torque $T_{e,i}$, which can be expressed as:
$T_{e}=T_{e,d}+T_{e,i}$ (17)
[…]
where $T_{e,max}=f_{e}(\omega_{e})$ represents the maximum engine torque at a given angular velocity $\omega_{e}$, which is depicted in Fig. 3. When the clutch is disengaged, $T_{e,d}$ is set to zero, and $T_{e}=T_{e,i}$. The idle torque primarily serves to power the auxiliary systems. The engine speed is governed by the following equation:
$\omega_{e}=\left\{\begin{array}[]{l}\omega,\quad\text{if }c=1\\ \omega_{ei},\quad\text{if }c=0\\ \end{array}\right.$ (21)
[…]
where $\omega_{m}$ is the motor angular velocity, and $\omega_{ei}$ is the engine idle angular velocity. The constraint on $\omega_{e}$ is imposed to ensure that the engine speed does not drop below the idle speed when the clutch is engaged.
The fuel rate can then be described by:
$\dot{m}_{f}=f_{f}(\omega_{e},T_{e})$ (23)
where $\dot{m}_{f}$ is the fuel rate (kg/s), and $f_{f}$ is the fuel rate interpolation function, which is shown in Fig. 3.
III-D Motor Model
Figure 4: The motor efficiency map, where the upper and lower boundaries represent the maximum and minimum motor torque, respectively.
The motor torque $T_{m}$ is governed by the following constraints:
$-f_{m}(\omega)=T_{m,min}\leq T_{m}\leq T_{m,max}=f_{m}(\omega)$ (24)
where $T_{m,max}$ and $T_{m,min}$ are the maximum and minimum motor torque at a given angular velocity, respectively. $f_{m}$ is an interpolation function that is used to calculate $T_{m,max}$ and $T_{m,min}$, which is depicted in Fig. 4.
III-E Battery Model
Figure 5: The open-circuit voltage and resistance of the battery cell.
We consider a lithium-ion battery pack consisting of 112 cells, where the cells are connected in series. The battery model is based on the battery state of charge (SOC) and the battery open-circuit voltage $E$. The battery SOC is calculated by:
$\dot{SOC}=-\frac{E-\sqrt{E^{2}-4R_{b}P_{b}}}{2R_{b}Q_{\max}}$ (25)
$P_{b}=T_{m}w\eta_{m}^{\operatorname{sgn}\left(-T_{m}\right)}$ (26)
$\eta_{m}=f_{\eta}(\omega,T_{m})$ (27)
where $I_{b}$ is the net current, $Q_{max}$ is the battery capacity, $R_{b}$ is the battery resistance, $P_{b}$ is the battery power, $\eta_{m}$ is the motor efficiency, and $f_{\eta}$ is the motor efficiency interpolation function, which is shown in Fig. 4. The battery open-circuit voltage and resistance are denoted as functions of the battery SOC, as illustrated in Fig. 5.
III-F Problem Formulation
The primary objective of the PHEV energy management system is to develop a control strategy that minimizes the equivalent cost of fuel and electricity consumption, while simultaneously reducing the frequency of gear shifts and clutch engagement to enhance the overall driving experience. The control inputs to the system are the engine torque, mechanical brake torque, gear ratio, and clutch state:
$u$ $=[T_{e},T_{b},\varsigma,c]^{\top}$ (28)
assuming that only a sequential shift pattern is considered, $\varsigma$ belongs to the set $\{-1,0,1\}$, where $\varsigma=-1$ represents downshift, $\varsigma=1$ represents upshift, and $\varsigma=0$ corresponds to sustainment. The cost function to minimize can be expressed as:
$J=\sum_{T_{0}}^{T_{cyc}}(k_{f}\dot{m}_{f}+k_{b}P_{b}+p_{\varsigma}+p_{c})\Delta t$ (29)
$p_{\varsigma}=\zeta\cdot\tilde{p},\quad\text{if }|gear_{t}-gear_{t-1}|>0$ (30)
$p_{c}=\xi\cdot\tilde{p},\quad\text{if }|c_{t}-c_{t-1}|>0$ (31)
where $k_{f}$ and $k_{b}$ are the fuel and electricity prices, respectively, $\tilde{p}$ is the reference penalty, which is set to be the intermediate value of fuel consumption, $p_{\varsigma}$ and $p_{c}$ are the penalties for gear shift and clutch engagement/disengagement, respectively, $\zeta$ and $\xi$ are the penalty coefficients, $\Delta t$ is the control interval, $T_{0}$ and $T_{cyc}$ are the initial and final time steps of the drive cycle, respectively. The drive cycle utilized for training is the China heavy-duty transient cycle for light trucks (CHTC-LT).
IV Experiments and Discussion
IV-A TD3AQ for PHEV Energy Management
IV-A1 State
The observation space of the RL agent is defined as the state of the system, which is a vector of the following variables:
$s=[v,a,T_{w},SOC,i_{g},c]$ (32)
IV-A2 Action
The action space is defined by the torque relationship in Eq. (18) as follows:
$a=[T_{e},\varsigma,c]$ (33)
where $T_{e}$ is continuous, and $\varsigma$ and $c$ are discrete, making it an MIOC problem. In negative demand torque, regenerative braking is prioritized for battery charging. If regenerative braking is insufficient, mechanical braking provides the remaining torque. In positive demand torque, the mechanical braking torque is set to zero, and the engine and motor supply the demand torque. Thus, the proposed action determines the vehicle’s torque distribution entirely.
To handle the multi-discrete action space involving two actions, $\varsigma$ and $c$, we employ a composite action vector. This vector, with a length of $3\times 2=6$, represents specific combinations of three gear shift actions and two clutch actions. This approach effectively manages the multi-discrete action space while maintaining a compact representation.
IV-A3 Reward
In addition to the cost function defined in Eq. (29), it is imperative to ensure that the system dynamics remain within specific constraints. Thus, the reward function is designed as a combination of fuel consumption and the penalty for violating the constraints. The detailed hand-crafted reward function is defined as follows:
$r=r_{cost}-p_{\omega}-p_{T_{m}}-p_{SOC}$ (34)
$r_{cost}=-(k_{f}\dot{m}_{f}+k_{b}P_{b}+p_{\varsigma}+p_{c})$ (35)
$p_{\omega}=\tilde{p},\quad\text{if }\omega>\omega_{max}$ (36)
$p_{T_{m}}=10\tilde{p},\quad\text{if }T_{m}>T_{m,max}$ (37)
$p_{SOC}=\begin{cases}0,&\text{if }SOC_{l}\leq SOC\leq SOC_{h}\\ 10\tilde{p}\cdot\frac{|SOC-SOC_{h}|}{1-SOC_{h}},&\text{if }SOC>SOC_{h}\\ 10\tilde{p}\cdot\frac{|SOC-SOC_{l}|}{SOC_{l}},&\text{if }SOC<SOC_{l}\end{cases}$ (38)
where $p_{\omega}$, $p_{T_{m}}$, and $p_{SOC}$ are the penalties for exceeding the constraints on the motor torque, shaft angular velocity, and SOC, respectively. The penalty for exceeding the SOC constraint is designed as a linear function of the SOC deviation from the desired range, which extends from the lower bound $SOC_{l}$ to the upper bound $SOC_{h}$. Therefore, the greater the deviation of the SOC from the desired range, the higher the penalty.
IV-B Comparison Benchmarks
Four benchmark methods, DP, DDAQ [25], PATD3 [14], and Rainbow [11], are selected for comparison. DP employs the MATLAB toolbox developed by Sundström et al. [28]. The continuous engine torque is discretized with an interval of 25 Nm. Empirical experiments have demonstrated that finer discretization does not improve the performance significantly (less than 0.05% improvement in cost with 10 Nm discretization while requiring seven times more computation time). Since Rainbow is designed for discrete action spaces, it adopts the same discretization scheme as DP. All RL algorithms utilize Optuna [29] to tune the hyperparameters.
IV-C Experimental Setup
The training process involves a total of 200 000 timesteps, where each timestep represents one second of driving time. Model evaluation was performed every 10 000 timesteps, with each evaluation period lasting one episode, i.e., one drive cycle. The final model for testing was selected based on the highest return achieved during evaluation. All experiments were conducted on a desktop computer equipped with an AMD Ryzen 9 5950X CPU and an NVIDIA GeForce RTX 3080 Ti GPU. The source code of the algorithm is available at https://github.com/xujinming01/TD3AQ.
IV-D Optimal Control Performance
(a) Engine torque
(b) Motor torque
(c) Gear shift sequence
(d) Clutch state (0 = disengaged, 1 = engaged)
Figure 6: The hybrid action output of DP and TD3AQ for CHTC-LT.
TABLE I: Simulation Results for CHTC-LT
Strategy $\tau_{sh}$ $\tau_{cl}$ $\nu_{T_{m}}$ $\nu_{\omega}$ $\nu_{soc}$ Cost/¥ Gap
DP 52 32 - - - 15.04 -
TD3AQ 116 56 0 0 0 15.66 4.13%
DDAQ 87 48 0 0 0 16.21 7.80%
PATD3 226 174 1 5 0 16.58 10.21%
Rainbow 251 208 0 1 0 16.22 7.85%
Fig. 6 displays the hybrid action output of DP and TD3AQ employed in CHTC-LT. It is evident that TD3AQ exhibits similar output trends to DP, however, with a higher frequency of gear shifts and clutch state changes. Notably, DP operates under the assumption that the entire drive cycle is known in advance. In contrast, RL algorithms only possess access to the current state, yielding suboptimal solutions. Table I presents the simulation results for CHTC-LT, where $\tau_{sh}$ and $\tau_{cl}$ denote the number of gear shifts and clutch engagements, respectively, $\nu_{T_{m}}$, $\nu_{\omega}$, and $\nu_{soc}$ denote the instances where motor torque, shaft angular velocity, and SOC exceed the predefined constraints. Among the RL algorithms, TD3AQ exhibits the lowest cost, with a gap of 4.13% compared to DP. Furthermore, TD3AQ and DDAQ do not violate any constraints, while PATD3 and Rainbow demonstrate a few violations in terms of motor torque and shaft angular velocity constraints.
Figure 7: SOC trajectories of CHTC-LT drive cycle.
Fig. 7 depicts the SOC trajectories for the RL algorithms and DP. It is observed that the RL algorithms generally exhibit similar SOC trajectories to DP, except during the final portions of the drive cycle. Given that the lower bound $SOC_{l}$ is set to 0.3, all the algorithms reach an SOC of 0.3 by the end of the drive cycle, with the exception of DDAQ and PATD3. Importantly, it is worth emphasizing that there are no violations of the SOC constraint, indicating that RL algorithms can effectively learn to regulate the SOC within the desired range without prior knowledge of the specific drive cycle.
TABLE II: Wall-Clock Time Comparison for CHTC-LT
Strategy Training time (h) step time (ms)
TD3AQ 0.18 0.38
DDAQ 0.17 0.36
PATD3 0.18 0.74
Rainbow 0.59 0.97
DP N/A 96599
Table II presents the time consumption comparison for CHTC-LT, where step time represents the time required to execute a single action during testing. While DP can obtain the optimal solution for each step in a single calculation, it comes at the expense of a significant amount of time, with its single solution time being six orders of magnitude greater than RL. This substantial discrepancy in time consumption highlights the efficiency of the RL algorithms, as indicated by their step times falling within the millisecond range. Consequently, RL demonstrates promising potential for real-time optimal control applications.
IV-E Generalization Results
TABLE III: Generalization Results of the New Cycles
Test cycle Strategy $\nu_{T_{m}}$ $\nu_{\omega}$ $\nu_{soc}$ Cost/¥ Gap
JE05 DP - - - 13.60 -
TD3AQ 1 0 0 14.04 3.23%
DDAQ 0 0 0 14.61 7.41%
PATD3 6 2 0 14.77 8.60%
Rainbow 0 0 0 14.76 8.48%
HD-UDDS DP - - - 9.27 -
TD3AQ 4 0 0 9.88 5.94%
DDAQ 0 8 0 10.55 13.18%
PATD3 14 7 0 9.97 6.95%
Rainbow 10 7 0 10.38 11.36%
WHVC DP - - - 5.51 -
TD3AQ 0 0 0 5.72 3.63%
DDAQ 0 0 0 5.89 6.78%
PATD3 0 1 0 5.80 5.16%
Rainbow 1 0 0 6.13 10.99%
To evaluate the generalization capabilities of the proposed CDRL algorithm TD3AQ, the model trained exclusively on the CHTC-LT drive cycle is tested on three additional drive cycles: JE05, HD-UDDS, and WHVC. The simulation results are summarized in Table III.
The findings demonstrate that the TD3AQ algorithm exhibits commendable generalization performance when applied to new drive cycles. In comparison with other RL algorithms, TD3AQ closely approximates the cost achieved by the DP strategy, with an average gap of 4.27% across the three cycles. However, the remaining RL algorithms display a larger deviation in terms of cost.
Furthermore, it is worth noting that all RL algorithms exhibit a higher frequency of constraint violations in the new drive cycles, particularly PATD3 and Rainbow. This issue primarily arises from the inherent limitation of RL in enforcing hard constraints during the training process.
V Conclusion
In this brief, we present a novel reinforcement learning algorithm called TD3AQ for solving mixed-integer optimal control problems. The TD3AQ algorithm uses actor-Q networks to directly generate continuous and discrete actions from the current state. We evaluate the proposed TD3AQ algorithm on a hybrid electric vehicle energy management problem and show that it achieves satisfactory performance with a cost gap of 4.13% compared to the DP strategy. TD3AQ’s ability to generalize is demonstrated through the results on three new drive cycles. The algorithm’s sub-millisecond response time indicates its significant potential for real-time, mixed-integer optimal control applications. However, note that the algorithm carries the risk of violating constraints, which, if unaddressed, could potentially lead to vehicle and battery damage in real-world scenarios. Future work will focus on extending TD3AQ to more complex, higher-dimensional problems to further substantiate its adaptability and robustness, and developing CDRL algorithms with safety guarantees.